OpenStudy (anonymous):

For three consecutive integers the sum of the first and twice the second is 23 more than twice the third. Which equation can be used to find the three integers?

OpenStudy (anonymous):

If you call them \(x\), \(y\), and \(z\), you know that: \[y = x + 1\] \[z = y + 1 = x + 2\] Then, you can plug those formulas into: \[x+2y = 23 + 2z\] Do you understand how I got all of those?